# Simple Yet Effective: Extracting Private Data Across Clients in Federated Fine-Tuning of Large Language Models

### v1

@@ Abstract
Federated fine-tuning of large language models (FedLLMs) presents a promising approach for achieving strong model performance while preserving data privacy in sensitive domains. However, the inherent memorization ability of LLMs makes them vulnerable to training data extraction attacks. To investigate this risk, we introduce simple yet effective extraction attack algorithms specifically designed for FedLLMs. In contrast to prior "verbatim" extraction attacks, which assume access to fragments from all training data, our approach operates under a more realistic threat model, where the attacker only has access to a single client’s data and aims to extract previously unseen personally identifiable information (PII) from other clients. This requires leveraging contextual prefixes held by the attacker to generalize across clients. To evaluate the effectiveness of our approaches, we propose two rigorous metrics—coverage rate and efficiency—and extend a real-world legal dataset with PII annotations aligned with CPIS, GDPR, and CCPA standards, achieving 89.9% human-verified precision. Experimental results show that our method can extract up to 56.57% of victim-exclusive PII, with "Address," "Birthday," and "Name" being the most vulnerable categories. Our findings underscore the pressing need for robust defense strategies and contribute a new benchmark and evaluation framework for future research in privacy-preserving federated learning.
@@ 1 Introduction
Fine-tuning large language models (LLMs) using federated learning (FL) algorithms (Ye et al., 2024a, b; Zhang et al., 2023; Chen et al., 2024; Yao et al., 2024) enables distributed training across clients without requiring centralized data. Recent advances (Bai et al., 2024; Wu et al., 2025; Li et al., 2020; Karimireddy et al., 2020) have focused on algorithmic enhancements to improve performance and reduce training costs. However, the threat of training data extraction from federated LLMs remains largely underexplored.
LLMs are known to memorize and regurgitate portions of their training data (Carlini et al., 2021, 2023), including sensitive content such as Personally Identifiable Information (PII) (Shao et al., 2024; Kim et al., 2023; Nakka et al., 2024). In FL, clients retain local datasets and collaboratively train models by sharing intermediate updates (e.g., checkpoints) rather than raw data. Prior attacks often rely on unrealistic assumptions, such as the adversary possessing large portions of the training corpus (Yu et al., 2023; Huang et al., 2022).
In this work, we consider a more realistic threat model in which a malicious client exploits contextual prefixes from its own data to extract previously unseen PII from other clients. For example, a prefix like "The trial of this case has now concluded. The plaintiff, (Name)..." in the attacker client can be used to prompt the global model and elicit sensitive information from other clients. Herein, we propose three simple yet effective attack strategies: (1) collect all possible contextual prefixes from the attacker dataset for querying federated LLMs, (2) sampling high-frequency prefixes from the attacker’s dataset to query the federated LLM, and (3) locally fine-tuning on prefix–PII pairs in a supervised manner to increase extraction success rates.
We evaluate our attack methods in a realistic setting by annotating a real-world legal dataset with PII labels aligned with privacy regulations such as CPIS, GDPR, and CCPA (see Acronyms List A). To assess attack effectiveness, we define two metrics: i) coverage, the proportion of target PII from other clients that can be recovered, and ii) efficiency, the amount of PII extracted within a limited query budget.
The experiments on our dataset show that the proposed extraction attacks can achieve up to 56.57% coverage, with "Address," "Birthday," and "Name" identified as the most vulnerable PII types. However, we observe diminishing returns in coverage as the prefix budget increases. Our enhancement strategies improve the diversity of extracted PII under fixed budget constraints. These findings underscore the tangible privacy risks in FedLLMs and emphasize the need for further research into effective defense mechanisms. In summary, our contributions are three-fold:
@@
    We propose three simple yet effective training data extraction attack methods for FedLLMs. Their effectivess is assessed with two evaluation metrics: coverage and efficiency. The new attack approaches are orthogonal to the existing attacking methods, such as gradient reconstruction and membership inference.
@@
    We conduct comprehensive experiments demonstrating that these attacks can achieve up to 56.57% coverage of victim-exclusive PII, effectively extracting diverse PII types. Our results also reveal a trade-off between coverage and extraction efficiency.
@@
    We develop a real-world benchmark dataset by extending a legal corpus with PII annotations aligned with regulatory standards (CPIS, GDPR, CCPA), thus addressing the scarcity of high-quality datasets for PII extraction research in federated learning.
@@ 2 Related Work
This study is related to the fields of data extraction attacks and federated learning. For the reader’s convenience, a brief introduction to these concepts is provided in Section B. In this section, we only review the work that is directly related to our method.
@@ 2.1 PII Extraction Attacks in LLM
A PII extraction attack broadly refers to any attack aimed at eliciting outputs from the model that contain PII (Lukas et al., 2023; Nakka et al., 2024; Huang et al., 2022). Both verbatim and associative techniques can be used to conduct such attacks.
@@ 2.2 Privacy Threats in Federated Learning
Threats in Federated Learning can be categorized into two main areas: security and privacy(Wang et al., 2024; Xie et al., 2024; Li et al., 2024a). Security threats typically aim to disrupt the entire FL system by invalidating model training(Shejwalkar and Houmansadr, 2021) and introducing backdoors(Bagdasaryan et al., 2020; Chang et al., 2024). In contrast, privacy threats have attracted more attention from researchers and focus on stealing confidential information from the FL system, such as inferring sensitive properties(Melis et al., 2019), reconstructing clients’ private datasets(Zhu et al., 2019; Geiping et al., 2020), and determining the membership and source of training data(Rashid et al., 2025; Vu et al., 2024; Hu et al., 2024). To achieve these attacks, researchers often make different assumptions regarding the attacker’s knowledge. Common assumptions typically fall into two dimensions: whether the attacker is a client or a serverChu et al. (2023), and whether the attacker is semi-honest(Applebaum, 2017; Hu et al., 2024) or malicious. These assumptions determine whether the attacker has access to gradients, local datasets, model parameters and the ability to manipulate them.
@@ 3.1 Data Sources and Preprocessing
The majority of our dataset comes from the Challenge of AI in Law (CAIL) (Li et al., 2024b), supplemented by smaller portions from CJRC (Duan et al., 2019) and JEC-QA (Zhong et al., 2020). CAIL is a renowned annual competition offering a variety of tasks in legal NLP. We specifically focus on two natural language generation (NLG) tasks-SUM and RC-and three NLP tasks-Match, Exam, and Cls. Detailed descriptions of each task can be found in Section D and Table 4 lists examples of each task.
Prior work (Zhang et al., 2023; Yue et al., 2023) has employed these datasets for LLM and FedLLM research. Based on their outputs, we further preprocess and curate the data, as described in Section E. Table 5 provides detailed dataset statistics.
@@ 3.2 PII Labeling
We reviewed the definitions and examples of PII in various legal provisions, including CPIS, GDPR, CCPA and Singapore PDPC (see Acronyms List A), and used them as references to establish a systematic PII labeling standard. We selected PII types relevant to the text modality, removing those related to the biological modality (e.g., facial recognition, fingerprint, DNA). We also removed PII types that are unlikely to appear in legal texts (e.g., browser history, SMS content, IP & MAC addresses), as well as those that are difficult to describe or evaluate (e.g., medical examination reports, psychological trends). Ultimately, we defined labeling guidelines encompassing 7 major categories and 36 subcategories. These standards are summarized in Table 6, and the distribution of labeled PII types is shown in Figure 1.
@@
We employed a combination of machine-assisted annotation and manual verification to label the data. For each major category of PII, we designed a specific prompt (as shown in Figure 5) and utilized GPT-4o (OpenAI et al., 2024) for initial labeling. Subsequently, we used Label Studio (Tkachenko et al., 2020-2025) and recruited five students with a foundational understanding of legal concepts and judgment skills to evaluate a subset of the annotated data. We selected 100 samples, which included over 6,000 PII labeling records, and evaluated them. The results of this evaluation are presented in Table 7.
@@ 4.1 Problem Definition
We investigate a novel class of data extraction attacks that is particularly realistic in the emerging federated large language model (FedLLM) setting. This attack fundamentally differs from traditional verbatim data extraction in its assumptions, setup, and goals.
First, unlike prior work where the attacker is assumed to have access to most of the training data, our setting assumes that the attacker possesses only a small subset of the overall training corpus. Second, in our formulation, the prefix and target suffix do not form a contiguous span within the original training data. Specifically, extraction prefixes are sampled from the attacker’s local dataset $D_{a}$, while the target suffixes exist solely in other clients’ data and are absent from $D_{a}$. As a result, each prefix must generalize across contexts to trigger the generation of previously unseen suffixes. Third, the attacker is not interested in extracting all training completions, but rather focuses on recovering specific, sensitive content, such as personally identifiable information (PII).
Formally, we consider an FL system comprising $c$ clients $\mathcal{C}=\{C_{1},C_{2},\dots,C_{c}\}$, where each client $C_{i}$ holds a local dataset $D_{i}$. Among them, one client—denoted as the attacker $C_{a}\in\mathcal{C}$—is assumed to be semi-honest (Applebaum, 2017; Hu et al., 2024). That is, $C_{a}$ faithfully follows the FL protocol (e.g., does not poison data or manipulate model weights), but acts adversarially in a passive manner, attempting to infer PII contained in other clients’ datasets by analyzing the global model $\theta$.
In this setting, the attacker issues queries to the model $\theta$ to extract data without knowing which client any particular output originates from. However, for evaluation purposes, we designate one client as the reference victim to measure the attack’s effectiveness. Let $S_{a}$ and $S_{v}$ denote the sets of PII instances held by the attacker and the victim client, respectively. The attacker constructs a prompt set $P$ and queries the federated language model $\theta$, obtaining a corresponding output set $Y$. We formalize key definitions and evaluation metrics in the following subsections.
@@
Definition 2 (Coverage Rate (CR)).
@@
$S_{E}$ =$=\{s_{i}\mid\exists y\in Y\text{ such that }s_{i}\text{ is extracted by }y\},$
CR $=\frac{|(S_{v}\setminus S_{a})\cap S_{E}|}{|S_{v}\setminus S_{a}|}.$ (2)
@@
Building upon these definitions, the central challenge is to design algorithms that enables the attacker to extract PII both comprehensively and efficiently—that is, achieving high coverage (CR) and high efficiency (EF).
@@ 4.2.1 PII-contextual Prefix Sampling
We begin with a straightforward method for constructing query prompts using PII-contextual prefixes—text segments that immediately precede PII instances in the attacker’s own dataset $D_{i}$. This approach is motivated by the observation that manually crafted prompts such as "my phone number is" may not align with the model’s training distribution, and are often ineffective at extracting attacker-defined PII instances.
Let the attacker’s training corpus be denoted as $U_{a}=t_{0}t_{1}\cdots t_{|U_{a}|}$, formed by concatenating all samples in $D_{i}$, where each $t_{i}$ represents a word token. Let $\mathcal{S}$ be the multiset of PII instances labeled in $U_{a}$. Define the function $\text{Loc}(s)$ as the index of the first word of a PII instance $s$ in $U_{a}$. We further define a contextual prefix extraction function $\mathcal{T}_{\lambda}(U,s)$ that returns the $\lambda$-length prefix ending just before $s$:
@@
The resulting set of PII-contextual prompts is given by:
@@
Once $P_{c}$ is constructed, the attacker $\mathcal{C}_{a}$ uses the global model $\theta$ to generate candidate PII completions. For each prefix $p\in P_{c}$, a suffix $y$ of at most $m$ tokens is sampled according to the model’s conditional distribution:
$y=x_{1}x_{2}\dots x_{m}\sim\mathbf{P}(y\mid p;\theta).$
To increase the diversity of generations, $\mathcal{C}_{a}$ may sample $n$ independent suffixes for each prefix $p$, forming:
$Y_{p}=\{y_{1},y_{2},\dots,y_{n}\}.$
The final set of candidate generations is the union across all prefixes:
$Y=\bigcup_{p\in P_{c}}Y_{p}.$
This results in a total query cost of $Q=n\cdot|P_{c}|$.
We can further generalize the contextual prompt set $P_{c}$ by collecting a multiset of all prefix substrings that end immediately before each PII instance:
@@
This produces a much larger set of contextual prefixes, which may yield a higher extraction coverage rate (CR) but an extremely low efficiency (EF) due to the massive number of query prompts.
4.2.2 Frequency-Prioritized Prefix Sampling (FP Sampling)
We hypothesize that prefixes appearing more frequently immediately before PII entities are more strongly associated with different PII instances. Based on this intuition, our goal is to prioritize such high-frequency prefixes to construct a more compact and informative prefix set.
To formalize this, we partition the multiset $\text{SUP}(P_{c})$ based on prefix frequency. For each integer $\sigma\geq 1$, let:

### v2

@@ Abstract
Federated large language models (FedLLMs) enable cross-silo collaborative training among institutions while preserving data locality, making them appealing for privacy-sensitive domains such as law, finance, and healthcare. However, the memorization behavior of LLMs can lead to privacy risks that may cause cross-client data leakage. In this work, we study the threat of cross-client data extraction, where a semi-honest participant attempts to recover personally identifiable information (PII) memorized from other clients’ data. We propose three simple yet effective extraction strategies that leverage contextual prefixes from the attacker’s local data, including frequency-based prefix sampling and local fine-tuning to amplify memorization. To evaluate these attacks, we construct a Chinese legal-domain dataset with fine-grained PII annotations consistent with CPIS, GDPR, and CCPA standards, and assess extraction performance using two metrics: coverage and efficiency. Experimental results show that our methods can recover up to 56.6% of victim-exclusive PII, where names, addresses, and birthdays are particularly vulnerable. These findings highlight concrete privacy risks in FedLLMs and establish a benchmark and evaluation framework for future research on privacy-preserving federated learning. Code and data are available at https://github.com/SMILELab-FL/FedPII.
Simple Yet Effective: Extracting Private Data Across Clients in Federated Fine-Tuning of Large Language Models
Yingqi Hu1, Zhuo Zhang1, Jingyuan Zhang2, Jinghua Wang1, Qifan Wang3, Lizhen Qu4thanks: Corresponding author., Zenglin Xu5,6 1Harbin Institute of Technology, Shenzhen, China 2Kuaishou Technology, China  3Meta AI, USA 4Monash University, Australia  5Fudan University, China 6Shanghai Academy of Artificial Intelligence for Science, China
@@ 1 Introduction
Federated large language models (FedLLMs) (Ye et al., 2024a, b; Zhang et al., 2023; Chen et al., 2024; Yao et al., 2024) have recently emerged as a promising approach for cross-silo federated learning (FL) (Li et al., 2024c), enabling collaborative model training while maintaining data locality and institutional privacy. In cross-silo FL, organizations, such as courts, banks, and hospitals, collaboratively fine-tune [^1] In current practice, FedLLM typically refers to the federated fine-tuning of large language models rather than federated pre-training. See Appendix B.3 for details. a shared model without exchanging private records. Prior studies mainly focused on improving algorithmic efficiency and convergence (Bai et al., 2024; Wu et al., 2025; Li et al., 2020), while the privacy vulnerabilities of FedLLMs remain largely unexplored.
A growing body of work has shown that large language models (LLMs) tend to memorize and reproduce fragments of their training data, including PII such as names, addresses, and dates of birth (Carlini et al., 2021, 2023; Shao et al., 2024; Kim et al., 2023; Nakka et al., 2024). Although FL mitigates privacy risks by exchanging model updates instead of raw data, our preliminary experiments (Appendix B.4) indicate that FedLLMs remain vulnerable to verbatim data extraction (VDE)—where adversaries can recover verbatim text sequences from the aggregated global model. However, most existing VDE studies assume that attackers have privileged access or significant knowledge of the victim’s data (Yu et al., 2023; Huang et al., 2022), which is unrealistic in practical cross-silo deployments.
Figure 1: Overview of cross-silo FedLLMs and the proposed privacy attack. In cross-silo FL, institutions such as banks, courts, and hospitals collaboratively fine-tune a shared model under the coordination of a central server, keeping data local. A semi-honest client leverages its local data to construct PII-related prefixes and queries the aggregated global FedLLM, leading to cross-client data leakage. The proposed strategies—Contextual Prefix Sampling, Frequency-prioritized (FP) Sampling, and Latent Association Fine-tuning (LAFt)—achieve up to 56.6% recovery of victim-exclusive PII, with names, addresses, and birthdays being the most vulnerable categories.
We instead consider a more realistic semi-honest threat model, where each participant follows the FL protocol but may attempt to infer private information from the global model. For instance, in a federation of courts, a participant could exploit its own local case records as contextual prefixes to elicit sensitive information memorized from other courts’ data (as illustrated in Figure 1). To investigate this, we propose three extraction strategies: (1) PII Contextual Prefix Sampling, which queries the global model using local contextual prefixes; (2) Frequency-prioritized (FP) Sampling, which focuses on high-frequency prefixes to improve extraction efficiency; and (3) Latent Association Fine-tuning (LAFt), which locally fine-tunes the global model using stored prefix-PII pairs, thereby improving its ability to extract PIIs implicitly memorized within the model.
To evaluate these attacks, we build a benchmark dataset by annotating a real-world Chinese legal corpus with fine-grained PII labels aligned with privacy regulations such as CPIS, GDPR, and CCPA (see Acronyms List A). We assess the attacks using two metrics: coverage (the proportion of target PII successfully extracted) and efficiency (the amount of PII recovered under a limited query budget). Our experiments reveal that the proposed attacks can extract up to 56.6% of victim-exclusive PII, with Name, Address, and Birthday being the most vulnerable types. Moreover, we observe diminishing returns as query budgets grow, while FP Sampling and LAFt enhance diversity under tighter budgets. These findings expose concrete privacy risks in FedLLMs and underscore the need for stronger privacy-preserving mechanisms.
In summary, our main contributions are:
@@
    We propose three novel extraction strategies for FedLLMs, independent of existing gradient-based or membership inference attacks, and evaluate them using two rigorous metrics: coverage and efficiency.
@@
    Extensive experiments show that our methods can recover up to 56.6% of cross-client unique PII, with larger prefix sets yielding diminishing returns in efficiency, revealing a trade-off between coverage and computational cost.
@@
    We construct a real-world benchmark dataset by augmenting a legal-domain corpus with fine-grained PII annotations aligned with CPIS, GDPR, and CCPA standards, filling the gap in public resources for privacy research in federated learning.
@@ 2 Related Work
This study is related to the fields of data extraction attacks and federated learning. For the reader’s convenience, a brief introduction to these concepts is provided in Appendix B. In this section, we review only the work directly related to our method.
@@ 2.1 PII Extraction Attacks in LLM
Broadly, PII extraction attacks refer to any attack that aims at eliciting outputs from the model that contain PII (Lukas et al., 2023; Nakka et al., 2024; Huang et al., 2022). Both verbatim and associative techniques can be used to conduct such attacks.
While most prior work assumes centralized training with full data access, we investigate PII extraction under federated fine-tuning, where the attacker has limited observability and control. We elaborate on this in Section 4.1.
@@ 2.2 Privacy Threats in Federated Learning
Threats in Federated Learning can be categorized into two main areas: security and privacy (Wang et al., 2024a; Xie et al., 2024; Li et al., 2024b). Security threats typically aim to disrupt the entire FL system by invalidating model training (Shejwalkar and Houmansadr, 2021) and introducing backdoors (Bagdasaryan et al., 2020; Chang et al., 2024). In contrast, privacy threats have attracted more attention from researchers and focus on stealing confidential information from the FL system, such as inferring sensitive properties (Melis et al., 2019), reconstructing clients’ private datasets (Zhu et al., 2019; Geiping et al., 2020), and determining the membership and source of training data (Rashid et al., 2025; Vu et al., 2024; Hu et al., 2024). To achieve these attacks, researchers often make different assumptions regarding the attacker’s knowledge. Common assumptions typically fall into two dimensions: whether the attacker is a client or a server (Chu et al., 2023), and whether the attacker is semi-honest (Applebaum, 2017; Hu et al., 2024) or malicious. These assumptions determine whether the attacker has access to gradients, local datasets, model parameters, and the ability to manipulate them.
@@ 3.1 Data Sources and Preprocessing
The majority of our dataset is sourced from the Challenge of AI in Law (CAIL) (Li et al., 2024a), supplemented by smaller portions from CJRC (Duan et al., 2019) and JEC-QA (Zhong et al., 2020). CAIL is a renowned annual competition featuring a variety of legal NLP tasks. In this study, we focus on two natural language generation tasks: Judicial Summary and Judicial Reading Comprehension, and three natural language understanding tasks: Similar Case Matching, Judicial Exam Question Answering, and Legal Case Classification. Detailed task descriptions are provided in Appendix D, with representative examples shown in Table 6.
Following prior work (Zhang et al., 2023; Yue et al., 2024), we further preprocess and curate the dataset to fit our setting. The complete preprocessing pipeline is described in Appendix E, where Table 7 reports the dataset statistics. [^2] The datasets contain PII from publicly available government-published legal documents. They were de-identified and used in prior work, e.g., as in Yue et al. (2024). We use curated versions from these papers. Since our study concerns privacy risks in FedLLMs, real-world PII is necessary to evaluate model vulnerabilities.
@@ 3.2 PII Labeling
We reviewed the definitions and examples of PII in various legal provisions, including CPIS, GDPR, CCPA, and Singapore PDPC (see Acronyms List in Appendix A), and used them as references to establish a systematic PII labeling standard. We selected PII types relevant to the text modality and removed types that are unlikely to appear in legal texts (e.g., browser history, SMS content, IP & MAC addresses), as well as those that are difficult to describe or evaluate (e.g., medical examination reports, psychological trends). Ultimately, we defined labeling guidelines encompassing 7 major categories and 36 subcategories. The distribution of labeled PII types is shown in Figure 2 in the main text, whereas a complete summary of these standards is provided in Table 9 in Appendix I.
We employed a combination of machine-assisted annotation and manual verification to label the data. For each major PII category, we designed a dedicated prompt depicted in Figure 12 and employed GPT-4o (OpenAI et al., 2024) to generate annotations. We then recruited students to verify a subset of those annotations with the help of Label Studio (Tkachenko et al., 2020-2025). The agreement between human evaluations and GPT-4o annotations achieves an F1 score of 89.9%. Owing to space limitations, the full evaluation results are shown in Table 10 in Appendix J, together with further details of the annotation process, including annotator backgrounds, annotation instructions, and user interfaces.
@@ 4.1 Problem Definition
We study a novel extraction attack tailored to FedLLMs, which differs from traditional verbatim data extraction in three key aspects:
Assumptions. Unlike VDE that assumes the attacker has access to most or all of the training data, our setting limits the attacker to a small, isolated subset of the overall training corpus.
Setup. In our formulation, the prefix and its corresponding target suffix are not drawn from a contiguous span of training data. Instead, extraction prefixes are sampled from the attacker’s local dataset $D_{a}$, while the target suffixes reside exclusively in other clients’ private data and are absent from $D_{a}$. Thus, each prefix must generalize beyond local context to trigger the generation of unseen suffixes.
Goals. An attacker does not aim to recover all training completions, but instead focuses on extracting specific, high-value information—most notably, PII—from the global model.
Formally, we consider a cross-silo FL system comprising $c$ clients $\mathcal{C}=\{C_{1},C_{2},\dots,C_{c}\}$, where each client $C_{i}$ holds a local dataset $D_{i}$. Among them, one client—denoted as the attacker $C_{a}\in\mathcal{C}$—is assumed to be semi-honest (Applebaum, 2017; Hu et al., 2024). That is, $C_{a}$ faithfully follows the FL protocol (e.g., does not poison data or manipulate model weights), but acts adversarially in a passive manner, attempting to infer PII contained in other clients’ datasets by analyzing the global model $\theta$.
In this setting, the attacker issues queries to the model $\theta$ to extract data without knowing which client any particular output originates from. However, for evaluation purposes, we designate one client as the reference victim to measuring the attack’s effectiveness. Let $S_{a}$ and $S_{v}$ denote the sets of PII instances held by the attacker and the victim client, respectively. The attacker constructs a prompt set $P$ and queries the FedLLM $\theta$, obtaining a corresponding output set $Y$. We formalize key definitions and evaluation metrics in the following subsections.
@@
Definition 2 (Coverage Rate).
@@
$S_{E}=\{s_{i}\mid\exists y\in Y\text{ such that }s_{i}\text{ is extracted by }y\},$
$\text{CR}=\frac{|(S_{v}\setminus S_{a})\cap S_{E}|}{|S_{v}\setminus S_{a}|}.$ (2)
@@
Building upon these definitions, the central challenge is to design algorithms that enable the attacker to extract PII both comprehensively and efficiently—that is, achieving high coverage and high efficiency.
@@ 4.2.1 PII-contextual Prefix Sampling
We start with a simple method for constructing query prompts using contextual prefixes of PII, which are word sequences immediately preceding PII instances in the attacker’s dataset $D_{i}$. This mitigates reliance on manually crafted prompts (e.g., ‘my phone number is’), which often deviate from the model’s training distribution and exhibit limited empirical applicability.
Let the attacker’s training corpus be $U_{a}=\{t_{0},t_{1},\dots,t_{|U_{a}|}\}$, formed by concatenating samples in $D_{i}$, with $\mathcal{S}$ the multiset of labeled PII. For a PII instance $s\in\mathcal{S}$, let $\text{Loc}(s)$ be the index of its first token in $U_{a}$. We define a $\lambda$-length contextual prefix function:
@@
The contextual prefix set of a PII set $\mathcal{S}$ is given by:
@@
For each $p\in P_{c}$, the attacker $\mathcal{C}_{a}$ queries the global model $\theta$ to generate a suffix $y$ of up to $m$ tokens:
$y=\{x_{1},\dots,x_{m}\}\sim\mathbf{P}(y\mid p;\theta).$
To enhance diversity, $n$ independent suffixes may be sampled per prefix:
$Y_{p}=\{y_{1},\dots,y_{n}\},\quad Y=\bigcup_{p\in P_{c}}Y_{p},\quad Q=n\cdot|P_{c}|.$
A generalized version extends $P_{c}$ by including all substrings ending before each PII:
@@
This yields broader coverage but incurs high query cost due to the large prefix set.
4.2.2 Frequency-Prioritized Prefix Sampling
Following prior work (Shao et al., 2024), which associates extraction effectiveness with co-occurrence frequency, we posit that prefixes frequently occurring before PII entities tend to capture stronger and more diverse associations. Therefore, we emphasize high-frequency prefixes to construct a compact, information-rich set.
Formally, we partition $\text{SUP}(P_{c})$ by prefix frequency. For each $\sigma\geq 1$,